OpenStudy (gabylovesyou):

Which factor do u cancel from the numerator and denominator to simplify, x^2 + 3x - 4 / x^2 - 16 @hartnn @ganeshie8 @amistre64

OpenStudy (anonymous):

x^2 + 3x - 4 = (x+4)(x-1) and x^2 - 16 = (x+4)(x-4) Thus, your fraction is equal to (x+4)(x-1)/(x+4)(x-4). Like terms cancel each other, right?

OpenStudy (gabylovesyou):

yes so u cancel x + 4

OpenStudy (anonymous):

yeah.

OpenStudy (gabylovesyou):

thank you :)

OpenStudy (anonymous):

You are welcome!